OpenStudy (anonymous):

PLEASE HELPP!!! what is the formula of a line that is perpendicular to y=1/3x+9 and includes the point (3,4)

OpenStudy (azureilai):

The slope of your second line is going to be the opposite reciprocal of your first line. The slope formula is y=mx+b in order to find b, plug in your slope and points into the formula (y-y1)=m(x-x1)

OpenStudy (anonymous):

please show work

OpenStudy (anonymous):

i dont understand how to do this

OpenStudy (azureilai):

Well, do you know what an opposite reciprocal is?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

@Azureilai ???

OpenStudy (anonymous):

@Azureilai hello?????

OpenStudy (azureilai):

Ok, so lets consider the fraction 1/2. the opposite reciprocal would be -2/1. The slope of a line is usually in a fraction, so what you do is you flip the numerator and the denominator of the fraction and negate it. (If the fraction was positive, it becomes negative, if it was negative then it becomes positive.) Can you follow so far?

OpenStudy (anonymous):

yes

OpenStudy (azureilai):

Good, so what do you think will be the opposite reciprocal of 1/3 ?

OpenStudy (anonymous):

-3/1

OpenStudy (azureilai):

yes you are correct. Now are you familiar with the point slope form?

OpenStudy (anonymous):

y=mx+b??

OpenStudy (azureilai):

Thst is the slope intercept form. Point slope form is (y-y1)=m(x-x1) You can use it to solve for b in the y=mx+b (since we have already determined the slope, which is -3, all we need now is the b value.) What you do in the new formula is plug in you points (3,4,) into the equation and isolate y again to one side. so (y-4)=-3(x-3) Solve for y

OpenStudy (anonymous):

@Azureilai i dont know how to solve for y

OpenStudy (azureilai):

Ok, lets consider the right part of the equation first, can you distribute the -3 in -3(x-3)?

OpenStudy (anonymous):

@Azureilai can u please draw it?

OpenStudy (anonymous):



OpenStudy (azureilai):

 do you get it?